OpenStudy (anonymous):

find domain of: 1/ [(the 4th root of)x^2 -9x]

OpenStudy (precal):

set the denominator equal to zero and solve it. Recall division can not have a zero, so the domain is all of the real numbers except those that make it zero.

OpenStudy (anonymous):

okay..does the 4th root change anything? thats what is confusing me

OpenStudy (anonymous):

\[x \neq 0\]\[x \neq 9\]

jhonyy9 (jhonyy9):

no,the 4th root not change nothing

OpenStudy (anonymous):

ok

OpenStudy (precal):

I meant to say the denominator can not be equal to zero and division by zero is undefined.

OpenStudy (anonymous):

so i would set: \[\sqrt[4]{x^2 -9x}\] =0?

OpenStudy (anonymous):

and solve for that

OpenStudy (precal):

I meant to say the denominator can not be equal to zero and division by zero is undefined.

OpenStudy (precal):

Yes take both sides to the 4th power and that would get rid of the (1/4) power.

OpenStudy (anonymous):

\[x^2-9x \neq0\]

OpenStudy (anonymous):

so x^2 - 9x cant equal zero...so it equals zero at....\[\sqrt{9x}\] ? haha

OpenStudy (precal):

I meant to say the denominator can not be equal to zero and division by zero is undefined.

OpenStudy (precal):

factor \[x^2-9x\]

OpenStudy (anonymous):

ok

OpenStudy (precal):

I meant to say the denominator can not be equal to zero and division by zero is undefined.

OpenStudy (precal):

x(x-9) therefore x can not be zero and x can not be 9 but x can be any other real number.

OpenStudy (anonymous):

i see...so i didnt need to solve for x, just factor? to get x(x-9)

OpenStudy (anonymous):

\[x(x-9)\neq0 \rightarrow x \neq 9, x \neq 0\]

OpenStudy (precal):

I meant to say the denominator can not be equal to zero and division by zero is undefined.

OpenStudy (precal):

You are solving for zero.

myininaya (myininaya):

the domain is \[x^2-9x>0\]

OpenStudy (anonymous):

okay...i think i can take it from here..hopefully haha, thanks guys

OpenStudy (precal):

I meant to say the denominator can not be equal to zero and division by zero is undefined.

OpenStudy (precal):

what about x=9, x can not be 9, it would create a zero in the denominator. You can not have a zero in the denominator.

OpenStudy (anonymous):

The domain is \(x^2-9x \ge 0\).

myininaya (myininaya):

right thats why we have > and not >=

myininaya (myininaya):

anwar we don't want the bottom to be zero

myininaya (myininaya):

\[x^2-9x>0\]

OpenStudy (anonymous):

\[x <0, x >9\]

OpenStudy (anonymous):

how would i express x cant equal 0 and 9 in interval notation...im having trouble with that

OpenStudy (anonymous):

Oh it's in the denominator, I didn't see that :)

OpenStudy (anonymous):

you express it as \[(-\infty,0)\cup (9,\infty)\]

OpenStudy (anonymous):

hello all!

OpenStudy (anonymous):

ahh...i see where i made my mistake entering it, thank you!

OpenStudy (anonymous):

And hello to @myininaya and @satellite73 :)

myininaya (myininaya):

x(x-9)>0 --------------------------- x(x-9)=0 when x=0 or x=9 f(-1)=+ f(1)=- f(10)=+ so the intervals for which the function exist are \[(-\infty, 0) U (9,\infty)\] oops what satellite has lol

myininaya (myininaya):

hey anwar! :)

OpenStudy (anonymous):

lol @myininany \cup

myininaya (myininaya):

shhh

OpenStudy (anonymous):

@satellite73: You should include the real numbers between \(0\) and \(9\).

myininaya (myininaya):

no he shouldn't

OpenStudy (anonymous):

well not really since you cannot take the fourth root of a negative number

OpenStudy (anonymous):

@Anwara NO

OpenStudy (anonymous):

Oh no, these are the negative values, my bad!

myininaya (myininaya):

lol

OpenStudy (anonymous):

ok ;)

OpenStudy (anonymous):

haha..thanks guys i appreciate the help!

OpenStudy (anonymous):

you're welcome ;)